OpenStudy (anonymous):

Simplify log(base a)1-log(base a)a^b

OpenStudy (anonymous):

log(1/(a^b)) => log(a^(-b)) => -blog(a) => -b sing base is a we have -b

OpenStudy (anonymous):

Remember that by definition, for any number x, $$\log_x 1= 0,$$ since any number to the zeroth power is 1. Also, remember that, by definition for any number x, $$\log_x x^b = b.$$ That's what a logarithm is. So combining those two: $$\log_a 1 - \log_a a^b = 1 - b$$ That's right, the a doesn't matter at all!

OpenStudy (anonymous):

Sorry, I that last equation was supposed to read: $$\log_a 1 - \log_a a^b = -b$$

OpenStudy (anonymous):

Rule #1 of posting corrections: your correction will itself contain a mistake. <goes away grumbling darkly/>

OpenStudy (anonymous):

thanks, i couldnt put the equation thing. so the answer is just -b??? that simple?

OpenStudy (amistre64):

yes, -b

OpenStudy (anonymous):

but you said logx1=0 so shoudltn it be 0-b?

OpenStudy (amistre64):

0 is not a "simplification"; it is an overcomplication :)

OpenStudy (amistre64):

0 -b = -b ; see :)

OpenStudy (anonymous):

ok :) thanks